OpenStudy (anonymous):

(e^x/3 cos x)(e^x sin 6x) how about this multiplication of it?thanks in advance

OpenStudy (anonymous):

\[e ^{\frac{x^2}{3}}\cos(6x)\sin(6x)\]

OpenStudy (anonymous):

can u show me the full solution?i couldnt understand.

OpenStudy (anonymous):

well to figure out the e term, you use the rule \[(a^b)a^c= a^{b+c}\] so if you have\[e^{\frac{x}{3}}*e^x\] what do you get? btw, I actually got my e term wrong, so its not what i wrote before

OpenStudy (anonymous):

e^2/3 x is it correct??

OpenStudy (anonymous):

its actually \[e^{4x/3}\], because \[e^x = e^{3x/3}\]

OpenStudy (anonymous):

ow i see

OpenStudy (anonymous):

and because everything is timesed, you just take the e term you found out, and times it by the sin term and the cos term

OpenStudy (anonymous):

so you should get \[e^{\frac{4x}{3}}\cos(6x)\sin(6x)\]